OpenStudy (anonymous):

If F(x) = integral evaluated from 0 to x (sqrt((t^5)+4)dt) then find F'(2). Calc geniuses, help please!

OpenStudy (anonymous):

\[F(x)=\int_0^x(\sqrt{t^5}+4)dt\] like that?

OpenStudy (anonymous):

or \[F(x)=\int_0^x\sqrt{t^5+2}dt\] like that

OpenStudy (anonymous):

extend the sqrt over the four and you got it!

OpenStudy (anonymous):

this one \[F(x)=\int_0^x\sqrt{t^5+4}dt\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you are going to be disgusted at how easy this is

OpenStudy (anonymous):

by the fundamental theorem of calculus, if \[F(x)=\int_0^x\sqrt{t^5+4}dt\] then \[F'(x)=\sqrt{x^5+4}\]

OpenStudy (anonymous):

Wow. Well thank you:) and yes all the disgust right now haha

OpenStudy (anonymous):

and therefore \[F'(2)=\sqrt{2^5+4}\] whatever that is

OpenStudy (anonymous):

yw